OpenStudy (gabbyalicorn):

Could you double check these two questions please? @563blackghost Thanks Larry’s Landscaping charges $235 for spring cleanups and $30 for weekly lawn maintenance. Joe’s Landscaping charges $185 for spring cleanups and $35 for weekly lawn maintenance. The system that models this situation is given, where c is the cost of lawn maintenance and w is the number of weeks. c = 235 + 30w c = 185 + 35w The solution to the system is (10, 535). Which interpretation correctly describes the solution to the system of equations? A. Larry’s Landscaping will cha

OpenStudy (gabbyalicorn):

Larry’s Landscaping charges $235 for spring cleanups and $30 for weekly lawn maintenance. Joe’s Landscaping charges $185 for spring cleanups and $35 for weekly lawn maintenance. The system that models this situation is given, where c is the cost of lawn maintenance and w is the number of weeks. c = 235 + 30w c = 185 + 35w The solution to the system is (10, 535). Which interpretation correctly describes the solution to the system of equations? A. Larry’s Landscaping will charge more money on the tenth week by charging $535. B. Joe’s Landscaping will charge more money on the tenth week by charging $535. C. The cost for lawn maintenance is the same, $535, for both landscaping companies after 10 wk. D. A customer can have his or her lawn maintained for a maximum of 10 wk. The customer will pay a total of $535. I think B

OpenStudy (gabbyalicorn):

And last one Mr. and Mrs. Miller both get in their cars and begin to drive at the same time. Mr. Miller is 280 mi from their house and drives directly home at a constant rate of 40 mph. Mrs. Miller is at their house and drives away from the house at a constant rate of 30 mph. Let d represent distance from home, and let t represent time in hours. Which system models this situation? A. d = 280 + 40t d = 30t B. d = 280 − 30t d = 40t C. d = 280 − 40t d = 30t D. d = 280 − 40t d = 280 + 30t Someone told me D and i just want to double check

OpenStudy (gabbyalicorn):

@563blackghost take your time :)

563blackghost (563blackghost):

The first question is incorrect.... To find the total amount we would input 10 weeks since we want to know how much cash in 10 weeks so we input it where w is.... what were your totals...?

OpenStudy (gabbyalicorn):

My totals were 350 for joe and 300 for larry 350>300

563blackghost (563blackghost):

hmmmm.... you did your calculations wrong... Larry's Calculations \[235+30(10)\] \[235+300\] \[=535\] Joe's Calculations \[185+35(10)\] \[185+350\] \[=535\] The totals should be equal...

OpenStudy (gabbyalicorn):

Ohhhhh.....My mistake...D:

563blackghost (563blackghost):

It's alright I took so long cause I messed up myself ^^

563blackghost (563blackghost):

im still checking for the second question...

OpenStudy (gabbyalicorn):

Okay, no worries! Take your time.

563blackghost (563blackghost):

I don't think its D... Mr. Miller is traveling back to his house meaning he is reducing his numbers so - and he travels at a rate of 40mph so that would put the equation as 280 - 40t... Now Mrs. Miller is "at" the house is traveling away from it at 30mph if we use the equation in D which is 280 + 30t we actually don't know how far she is traveling so it shouldn't be D so Mrs.Miller's equation should actually be set up as 30t....

563blackghost (563blackghost):

Do u get it?

OpenStudy (gabbyalicorn):

Yes, I do. Thank you very much. ^_^

563blackghost (563blackghost):

Np ^^ Glad I was of help!

OpenStudy (gabbyalicorn):

:)